a) graph f
b) compare the graph of f with y=x^2
How would I solve f(x)=2(x-1)^2-3?




Answer by rfer(16322)

2(X-1)(X-1)=
2(X^2-X-X+1)=
2X^2-4X+2-3=
2X^2-4X-1